### v1

# Is Translation All You Need? A Study on Solving Multilingual Tasks with Large Language Models

Abstract
Large language models (LLMs) have demonstrated strong multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora. Existing works leverage this phenomenon to improve their multilingual performances on NLP tasks. In this work, we extend the evaluation from NLP tasks to real user queries. We find that even though translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios. For culture-related tasks that need deep language understanding, prompting in the native language proves to be more promising since it can capture the nuances related to culture and language. Therefore, we advocate for more efforts towards the development of strong multilingual LLMs instead of just English-centric LLMs.
Is Translation All You Need?
A Study on Solving Multilingual Tasks with Large Language Models
[…]
1 Introduction
Large language models (LLMs) frequently demonstrate the capability to understand and generate text across multiple languages, a skill attributed to their training on vast corpora composed of texts from various languages OpenAI (2023); Shi et al. (2022); Muennighoff et al. (2023); Jiang et al. (2023); Nguyen et al. (2023). However, these datasets are often disproportionately dominated by English content Brown et al. (2020); Chowdhery et al. (2022); Workshop et al. (2023), resulting in an English-centric bias in LLMs. This imbalance can subsequently hinder the models’ proficiency in other languages, often leading to suboptimal performance in non-English contexts Ahuja et al. (2023); Lai et al. (2023); Zhang et al. (2023b).
Figure 1: Illustration of two kinds of LLMs on different tasks. "English-centric LLMs" refers to LLMs trained mainly in English corpora. "Multilingual LLMs" refers to ideal LLMs that are equally capable in all languages.
To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al. (2018); Ponti et al. (2020). Recent investigations have expanded this concept by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al. (2023); Qin et al. (2023); Etxaniz et al. (2023) for multilingual NLP tasks. For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a). Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then proceed to solve problems step by step, leads to better performance.
Figure 2: Examples illustrating how translation can both improve (top) and degrade (bottom) the performance of LLMs. The example in Chinese is drawn from MGSM and the example in Swahili is drawn from M3Exam, respectively. Translation is beneficial when the questions are language-independent. However, for questions requiring deep cultural knowledge, translation can hinder the ability to answer correctly.
Despite these advancements, the methodologies employed across different studies vary considerably, and the specific impact of translation in improving multilingual task performance has not been rigorously analyzed. Furthermore, the focus of these studies tends to be on specific NLP tasks and English-centric LLMs, with less attention given to real-world scenarios of multilingual user requests. This gap highlights a need for more nuanced research into the effectiveness of translation techniques across a spectrum of multilingual contexts.
In this study, we conduct a thorough investigation into the utility of translation with large language models for a variety of scenarios, including multilingual NLP tasks and real-world multilingual user queries. We compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages against several baselines, such as using native language prompts and cross-lingual prompts Huang et al. (2023). Our experiments also span across a diverse set of LLMs and translation tools to ensure the broad applicability of our findings.
Our results demonstrate that simply translating queries into English can already achieve the best results in multiple task categories. Subsequently, we investigate the utility of translation in real-world scenarios by conducting tests with requests from ShareGPT [^1] https://sharegpt.com/. We discover that translation tends to yield better responses for low-resource languages. However, when working with high-resource languages and advanced LLMs, employing prompts in native languages appears to be the more effective strategy. To further analyze such behaviors, we extend our evaluations to LLMs optimized for languages other than English, focusing specifically on culture-related tasks. We observe that for these models, prompts in the native languages yield superior results by capturing the nuances related to culture and language, as illustrated in Figure 2 (b). In summary, as illustrated in Figure 1, while translation can improve the performance of English-centric LLMs, it may not be optimal for all scenarios. In this work, we highlight the need for robust LLMs that are equally proficient across all languages to effectively address a broader range of cases. Therefore, we advocate for more efforts towards the development of multilingual LLMs.
2 Translation for NLP Tasks
In this section, we explore various prompting strategies across multiple languages and LLMs, encompassing a wide spectrum of NLP tasks. This allows us to understand how different prompting methods impact the performance of tasks that theoretically do not depend on any specific language.
2.1 Experiment Setup
Original Question
制作一件袍子需要 2 匹蓝色纤维布料和这个数量一半的白色纤维布料。它一共需要用掉多少匹布料
Native-Basic
{Original Question }
您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-Basic
{Original Question }
You should format your final answer as "Answer: <Arabic numeral>".
Native-CoT
{Original Question }
让我们一步步思考。
您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-CoT
{Original Question }
Let’s think step by step in English.
You should format your final answer as "Answer: <Arabic numeral>".
XLT
I want you to act as an arithmetic reasoning expert for Chinese.
Request: {Original Question}
You should retell the request in English.
You should do step-by-step answer to obtain a number answer.
You should step-by-step answer the request.
You should tell me the answer in this format ’Answer :’.
Trans-Google
Crafting a robe requires 2 bolts of blue fiber cloth and half that amount of white fiber cloth. How many pieces of fabric will it use in total?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
Trans-NLLB
To make a robe, two pieces of blue fiber and half of that amount of white fiber are needed. How many pieces of fabric does it take to make?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
Table 1: An example of zero-shot prompts for a Chinese problem. For Native-Basic, EN-Basic, Native-CoT, EN-CoT and XLT, we provide the original Chinese question as input and expect an answer in the corresponding format; for Trans-Google and Trans-NLLB, we input the translated question in English, and expect a step-by-step solution in English. To obtain the desirable output format, we instruct the models to output in specific format.
We perform assessments on six benchmarks on reasoning, understanding, and generation tasks that can encapsulate various abilities of LLMs. We follow the task and language selection from Huang et al. (2023) to consider a wide array of 24 diverse languages, including German (de), Russian (ru), French (fr), Chinese Simplified (zh), Spanish (es), Japanese (ja), Italian (it), Vietnamese (vi), Turkish (tr), Indonesian (id), Swahili (sw), Arabic (ar), Korean (ko), Greek (el), Thai (th), Bulgarian (bg), Hindi (hi), Estonian (et), Bengali (bn), Tamil (ta), Urdu (ur), Telugu (te), Haitian Creole (ht), and Southern Quechua (qu). We categorize languages larger than 1% frequency in Common Crawl [^2] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages as high-resource languages (i.e., de, ru, fr, zh, es, ja, it and vi), and the rest as low-resource languages. We exclude English since we want to evaluate the efficient prompting strategy for non-English tasks. As our prompting strategies involve methods with translators, we do not test on the FLORES-200 dataset.
2.1.1 Tasks
We evaluate the models and prompting strategies on the following tasks: 1) the MGSM benchmark Shi et al. (2022), which tests models on grade school math problems. 2) the XCOPA benchmark Ponti et al. (2020), which presents a premise with two choices to determine causality or effect; 3) the XNLI benchmark Conneau et al. (2018), which assesses Natural Language Inference by judging the relationship between a premise and a hypothesis; 4) the PAWS-X benchmark (Yang et al., 2019) involves identifying paraphrases in two sentences; 5) the MKQA dataset (Longpre et al., 2021), which focuses on open-domain questions; 6) the XL-Sum benchmark Hasan et al. (2021), which involves summarizing news articles. We choose a subset of 9 languages for MKQA and 5 languages for XL-Sum. For evaluation metrics across our study, we employ the token overlap F1 score specifically for the MKQA dataset, the ROUGE-1 score for assessing XL-Sum, and accuracy as the standard metric for all other benchmarks. More details of the benchmarks can be found in A.1.
For each task, we sample up to 500 examples from the test set for each language. For MGSM, since there are only 250 examples for each language, the entire test set will be used. For generation tasks like MKQA and XL-Sum, we will translate the answer back to the original language if the prompting strategy involves a translator.
2.1.2 Models
We conduct experiments with six LLMs, consisting of one closed-source language model and five open-source language models:
  • ChatGPT [^3] https://platform.openai.com/docs/models/gpt-3-5, which is the most capable and cost-effective model in the GPT-3.5 family and it has been optimized for chat. We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
  • Mistral-7B-Instruct (v0.2). This model is the instructed version of Mistral-7B Jiang et al. (2023).
  • Llama-2-13B-chat and Llama-2-70B-chat, which are the chat models in Llama-2 family Touvron et al. (2023). Due to computational resource limitations, we use the AWQ Lin et al. (2023) version for Llama-2-70B-chat.
  • bloomz-7b1, which is a model fine-tuned with multiple tasks, including multilingual tasks Muennighoff et al. (2023).
2.1.3 Prompting Strategies
Considering the language of instruction, whether to apply a chain-of-thought reasoning approach and the availability of translation tools, we assess the following experimental strategies in the experiment. We utilize a zero-shot approach across all strategies, given that the models we selected have been fine-tuned to follow instructions.
Basic prompt with native instructions (Native-Basic)
This is to directly pose the questions without employing additional prompting strategies such as chain-of-thought. Both the query information and the instruction provided are presented in the original language.
Basic prompt with English instructions (EN-Basic)
Compared with Native-Basic, EN-Basic  instructs LLMs with English but the query information is in the original language.
Native chain-of-thought (Native-CoT)
In Native-CoT, we ask the question in the native language and ask the model to reason with the native language with the instruction "Let’s think step by step." translated into that language.
English chain-of-thought (EN-CoT)
We pose the question in the native language but instruct the model to reason in English with the instruction "Let’s think step by step in English".
Cross-lingual-thought (XLT)
XLT Huang et al. (2023) is a state-of-the-art prompting method to handle multilingual NLP tasks. It prompts LLMs to translate the question into English and solve the problem step-by-step in English, leveraging the English-centric bias of LLMs.
Translate to English with Google Translate (Trans-Google)
It uses Google Translate API to translate the original questions into English and then solve the problem step by step.
Translate to English with NLLB models (Trans-NLLB)
Instead of using commercial translators, we use an open-source model, namely NLLB Team et al. (2022). Specifically, we chose nllb-200-3.3B to do the translation.
In addition to the prompting strategies, an output constraint is also included in the template to facilitate answer extraction. The settings are demonstrated in Table 1. In the case of open-source models, where the output format may deviate from the instructions, we utilize "Therefore, the answer <constraint> is" in appropriate languages in the second round to retrieve the ultimate answer.
2.2 Results and Discussions
Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
ChatGPT Native-Basic 34.4 72.3 52.2 49.7 35.4 31.9 46.0
EN-Basic 41.1 76.1 63.0 62.0 36.5 32.7 51.9
Native-CoT 49.9 72.4 50.6 50.3 35.4 31.1 48.3
EN-CoT 61.1 78.6 56.4 61.6 42.9 29.7 55.1
XLT 62.3 79.3 59.2 59.4 37.6 23.4 53.5
Trans-Google 73.9 84.5 60.5 67.2 43.8 30.3 60.0
Trans-NLLB 61.0 79.7 59.2 67.5 37.2 28.3 55.5
bloomz-7b1 Native-Basic 1.3 21.4 8.3 - 7.8 - 9.7
EN-Basic 2.0 56.5 43.9 - 10.6 - 28.3
Native-CoT 1.2 20.9 8.2 - 6.5 - 9.2
EN-CoT 1.7 53.9 35.9 - 9.3 - 25.2
XLT 1.7 50.5 35.4 - 8.0 - 23.9
Trans-Google 2.7 63.7 44.3 - 17.2 - 32.0
Trans-NLLB 2.4 61.8 43.8 - 14.7 - 30.7
Mistral-7B-Instruct Native-Basic 11.2 54.2 42.8 44.6 7.8 25.9 31.1
EN-Basic 23.8 34.9 50.2 46.9 7.0 21.6 30.7
Native-CoT 17.0 53.8 43.4 44.3 7.8 24.6 31.8
EN-CoT 27.6 40.8 50.0 46.6 11.5 18.7 32.5
XLT 31.8 61.5 46.0 47.8 9.6 17.6 35.7
Trans-Google 41.3 59.2 55.0 51.5 17.0 32.8 42.8
Trans-NLLB 31.7 54.4 53.0 52.4 15.5 32.2 39.9
Llama-2-13b-Chat Native-Basic 15.6 50.1 36.4 54.0 18.3 27.5 33.6
EN-Basic 19.0 54.3 43.4 59.1 20.2 31.1 37.9
Native-CoT 18.1 50.9 35.7 54.8 15.7 27.2 33.7
EN-CoT 19.9 54.5 43.7 57.6 19.8 30.2 37.6
XLT 22.3 56.0 51.4 55.7 19.0 21.2 37.6
Trans-Google 39.1 71.9 46.1 58.4 33.8 35.7 47.5
Trans-NLLB 31.8 68.2 45.4 57.8 28.4 34.2 44.3
Llama-2-70B-Chat Native-Basic 23.6 51.5 39.0 52.8 24.8 29.4 36.2
EN-Basic 28.6 55.3 46.5 60.4 24.7 31.2 40.1
Native-CoT 23.5 50.5 37.9 54.9 21.9 25.8 34.5
EN-CoT 30.2 61.2 45.9 64.9 31.1 30.2 42.3
XLT 32.8 58.7 52.2 55.7 26.6 17.7 38.4
Trans-Google 53.3 80.9 54.0 68.5 39.7 33.8 51.8
Trans-NLLB 43.8 77.1 52.2 69.2 19.4 33.1 49.1
Table 2: Average scores across different languages (excluding English) for the six benchmarks in zero-shot setting. The results of PAWS-X and XL-Sum for bloomz-7b1 are not considered since it was already pre-trained on these tasks. The best result for each model is in bold.
The main results are shown in Table 2. We also calculate the average performance for high-resource and low-resource languages separately for each task in Table 4 in Appendix A.2. We notice that Trans-Google, despite simple, surprisingly demonstrates the highest overall performance across various models and tasks. While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages. Besides this, we can have the following observations: 1) Trans-NLLB ranks second on average, trailing behind Trans-Google. The discrepancy between the two methods lies solely in the quality of translation, indicating that the NLLB model’s translation proficiency is inferior. Nonetheless, its open-source nature renders it a potentially favourable choice under specific circumstances. 2) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought. This finding aligns with those reported by Lai et al., 2023. 3) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM. For weaker models and tasks that can be answered directly, the basic prompt may be a better option. More results and discussions are in Appendix A.2.
3 Translation for Real User Queries
The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages. However, these benchmarks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries that often span an extensive range of topics and require nuanced comprehension. Moreover, these benchmarks are often constructed by translating the original English data into other languages Shi et al. (2022); Ponti et al. (2020); Conneau et al. (2018); Yang et al. (2019); Hasan et al. (2021). This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages.
To assess the impact of translation on real-world queries, we extract user requests from ShareGPT datasets. We select 10 languages ranging from high to low resources and randomly sample 50 requests for each language from the first round of the dataset. To evaluate the quality of the responses, we use GPT-4-Turbo (gpt-4-turbo-0125) as the judge. The prompt for the judge is shown in Figure 6 in the Appendix, which is adapted from Zheng et al. (2023). With this prompt, each response will get a score from 1 to 10. Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API. For the second option, we will translate the output back to the original language for consistency. We select the two best models on NLP tasks for evaluation: ChatGPT and Llama-2-70B-Chat.
3.1 Results and Discussions
In our comparison, we analyzed the scores of two sets of responses generated by the same model, calculating the win rate for each language. The results are shown in Figure 3 and we make the following observations: 1) ChatGPT shows different behaviours for high-resource languages and low-resource languages. For Japanese, Chinese, and Spanish, prompting with original queries has a higher win rate. For lower-resource languages, translation is often a better option. 2) For Llama-2-70B-Chat, translation has a higher win rate for all the languages. This echoes the fact that Llama-2-70B-Chat is an English-centric LLM. Even though translation may lose some information, the gain of better understanding after translation dominates the performance.
The distinct behaviours exhibited by Llama-2-70B-Chat and ChatGPT underscore the inherent differences between the LLMs. Llama-2-70B-Chat is predominantly English-centric, as evidenced by its improved performance in both NLP tasks and real user queries when input is translated into English. Conversely, ChatGPT demonstrates characteristics of a “multilingual LLM”, as shown in Figure 1(b), showcasing its broader linguistic capacity. Nevertheless, ChatGPT still relies on translation to bolster its performance for lower-resource languages, indicating room for improvement in its approach to truly multilingual processing.
To ascertain the extent to which answering user queries necessitates an understanding of local cultural nuances, we employed GPT-4-Turbo, utilizing a specially crafted prompt to analyze user queries across multiple languages, as shown in Figure 7 in the Appendix. The results, illustrated in Figure 4, reveal that a significant portion of the queries—ranging from 18% to 50% for each language—require knowledge of local culture for accurate responses. This variation underscores the rich cultural elements embedded within the data, affirming our hypothesis that queries containing culture-specific information pose a greater challenge for the LLMs.
(a) Win rate with ChatGPT
(b) Win rate with Llama-2-70B-Chat
Figure 3: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat.
3.2 Analysis
Figure 4: The percentage of the questions that necessitate local cultural knowledge.
ChatGPT has shown promising performance in some high-resource languages. We extend the experiment to investigate the effectiveness of multilingual LLMs in handling culture-specific tasks, aiming to assess how well these models grasp cultural nuances and idiomatic language. Through our experiment, we seek insights into the adaptability of LLMs across diverse cultural contexts.
3.2.1 Experiment Setup
We select the following models to represent various categories of LLMs and assess the potential differences in their response to translation:
  • ChatGPT and llama-2-70B-chat, which are the strongest models in our previous experiments.
  • Qwen1.5-72B-Chat Bai et al. (2023) and Yi-34B-Chat [^4] https://github.com/01-ai/Yi. These two open-source models demonstrate strong English and Chinese capabilities. We aim to test if they excel more in Chinese than in translating Chinese into English.
Please note that except ChatGPT, the results for the other three models are obtained by calling API from Together [^5] https://api.together.xyz/. We chose open-source models which are larger than 30B since we hypothesise that larger models contain more knowledge.
For the evaluation dataset, we choose M3Exam Zhang et al. (2023a), as the questions require strong multilingual proficiency and cultural knowledge to perform well. For the subject categories, we choose language and social science, which are more likely to contain language-dependent knowledge. For each language, we select up to 500 samples for evaluation. For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
3.2.2 Results and Discussions
Figure 5: Results of four LLMs on M3Exam (a) language and (b) social science subject categories. Note that in M3Exam, not all subjects are available in every language, resulting in a discrepancy in the number of languages for the two subjects.
Based on the results shown in Figure 5, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance for all languages. This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a. The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension. 2) Translation helps Llama-2-70B-chat in all the languages, suggesting that the model’s underperformance is due to poor language understanding rather than limitations of cultural knowledge. 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency. The translation process hurts Chinese performance, highlighting the significant influence of translation errors on comprehension. Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat.
From this experiment, it is evident that while translation serves as a straightforward and initially effective strategy to enhance the performance of English-centric LLMs on non-English tasks, it falls short of being optimal. Directly training LLMs in specific languages can yield superior results by eliminating the potential for translation errors.
4 General Discussions
In this work, we categorize multilingual tasks into two groups: language-independent tasks and language-dependent tasks. Among the NLP benchmarks discussed in Section 2, such as MGSM, XCOPA, XNLI, PAWS-X, and MKQA, are parallel benchmarks. This means that the examples maintain semantic equivalence across all languages, categorizing them as language-independent tasks. Language-independent tasks are typically easier to evaluate across different linguistic groups, as they rely on universal cognitive abilities, like reasoning and comprehension, rather than specific linguistic knowledge. In contrast, tasks involving real user queries and culture-related questions, discussed in Section 3, necessitate models to possess cultural and contextual understanding, categorizing them as language-dependent tasks. By distinguishing between language-dependent and language-independent tasks, we aim to promote fair and accurate evaluations across diverse linguistic contexts, enhancing the development and assessment of NLP systems in multilingual environments.
Corresponding to the two types of tasks, multilingual LLMs must develop dual competencies:
General Language Skills:
Essential for understanding and engaging with both task types, these skills enable LLMs to process and comprehend content across multiple languages, facilitating basic instruction interpretation and response generation.
Specific Language Skills:
These skills are crucial for tasks demanding an understanding of cultural nuances and linguistic subtleties, such as interpreting culturally specific queries.
By equipping LLMs with these comprehensive skills, we ensure their effectiveness in a wide array of multilingual contexts, enhancing both linguistic adaptability and cultural sensitivity.
The benefits of translation on NLP tasks demonstrate the dominance of English-centric LLMs. Even top models like ChatGPT, perform better through translation. In handling user queries, ChatGPT excels in native languages compared to translation, particularly in high-resource languages but struggles with low-resource ones. When testing on the M3Exam questions, Qwen1.5-72B-Chat and Yi-34B-Chat perform quite well in the native language for Chinese questions but have no such advantages for other languages. This dependency on translation highlights the challenges in developing LLMs that can equally handle a wide range of languages, especially those with limited resources. Such findings prompt further research and development efforts aimed at enhancing the multilingual efficacy of LLMs, striving for models that can understand and generate text across languages with minimal reliance on translation.
5 Conclusion
In this work, we perform a comprehensive assessment of LLMs across diverse multilingual tasks. These tasks include traditional NLP challenges, real user queries, and culture-specific tasks. We explore the use of translation-based methods as a strategy to overcome the limitations inherent in English-centric LLMs, thereby facilitating enhanced comprehension. However, we find that for tasks heavily reliant on language-specific nuances, employing prompts in the native language emerges as a more effective approach. This method is particularly adept at capturing the subtleties and intricacies unique to each language. Our findings suggest that the development of robust multilingual LLMs presents a more effective solution for addressing the challenges of multilingual tasks, especially for language-dependent tasks. Therefore, we advocate for the prioritization of research and development efforts towards the creation of strong multilingual large language models.
6 Limitations
This study aims to systematically assess the effectiveness of various prompting strategies across different tasks and LLMs. Due to limitations in computing resources, it was not possible to evaluate all existing prompting strategies comprehensively. However, we endeavoured to cover the most commonly employed strategies to formulate a broad conclusion. In our evaluation of LLMs on culture-related tasks, we specifically selected two LLMs optimized for Chinese, acknowledging it as one of the most widely spoken languages globally. The dataset used, M3Exam, comprises exclusively multiple-choice questions. It is important to note this specificity as it may influence the applicability of our findings. In our evaluation, we limited our sampling to up to 500 samples for each language within the benchmarks to manage computational constraints and ensure a broad yet feasible analysis scope. Consequently, our results might not be directly comparable with other studies that evaluate performance across the entire benchmark. In future work, we plan to extend our evaluation to LLMs optimized for other languages and to explore benchmarks presented in various formats beyond multiple-choice questions.
Acknowledgements
This research is supported, in part, by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore. Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and School of Computer Science and Engineering, Nanyang Technological University, Singapore, for their support.
Appendix A Additional Experiments
A.1 NLP Tasks
Here are the detailed descriptions of the NLP tasks:
Arithmetic Reasoning
The MGSM Shi et al. (2022) benchmark includes mathematical problems from grade school and requires the model to compute the accurate solution. It spans 10 languages, and we use the accuracy score for assessment.
Commonsense Reasoning
The XCOPA benchmark Ponti et al. (2020) consists of a single premise and two choices. The goal is to identify which choice is the cause or effect of the premise. It covers 11 languages from various families, with an accuracy score used for evaluation.
Natural Language Inference
The XNLI Conneau et al. (2018) benchmark includes one premise and one hypothesis. The model’s job is to determine if the hypothesis is entailed, contradicted, or neutral based on the premise. It covers 15 languages, and we evaluate it using the accuracy score.
Paraphrase Identification
The PAWS-X (Yang et al., 2019) benchmark consists of two sentences and requires the model to judge whether they are paraphrases. It covers 7 languages, and we assess based on accuracy score.
Question Answering
The MKQA dataset (Longpre et al., 2021) contains open-domain questions that require predicting short answers. Questions that are unanswerable or excessively long to have a specific answer are not considered during evaluation. This dataset covers 25 languages, with our focus on 9 languages: de, es, fr, ja, ru, th, tr, vi, and zh. We assess the model’s performance using the token overlap F1 score.
Summarization
The XL-Sum Hasan et al. (2021) benchmark requires the model to condense a lengthy news article into a brief summary. It covers 44 languages, and we select a subset of 5 languages: es, fr, tr, vi, and zh. We use the ROUGE-1 score for evaluation.
The prompts of EN-Basic for each task are shown in Table 3, which are adapted from Huang et al., 2023. The prompt templates for other prompting strategies and the instructions for output formats are designed according to the descriptions in Section 2.1.3.
A.2 Additional Results
The average performances for high-resource and low-resource languages are shown in Table 4. Table 5, Table 6, Table 7, Table 8, Table 9 and Table 10 shows the detailed results for MGSM, XCOPA, XNLI, PAWS-X, MKQA and XL-Sum, respectively. In addition to the finding in Section 2.2, We find XLT exhibits competitive performance in reasoning tasks; however, its performance in generation tasks is less impressive. Our findings indicate that when employing the XLT prompting strategy, ChatGPT declined to answer 26.4% of the questions in the XL-Sum tasks, responding with “I’m sorry, I cannot …” This refusal pattern was not observed when utilizing other prompting strategies. For open-source models, while we did not observe a refusal pattern, they do not follow the instructions properly, which also degrades their performance with XLT.
Benchmark #Test Basic Prompt
MGSM 250 {problem}
XCOPA 500
Here is a premise: {premise}. What is the {question}? Help me pick the more plausible option: -choice1: {choice1}, -choice2: {choice2}
XNLI 500
{premise} Based on previous passage, is it true that {hypothesis}? 1: Yes, 2: No, or 3: Maybe?
PAWS-X 500
Sentence 1: {sentence1} Sentence 2: {sentence2} Question: Does Sentence 1 paraphrase Sentence 2? 1: Yes, 2: No?
MKQA 500
Answer the question in one or a few words in {target_language}: {question}?
XL-Sum 500
{article} Summarize the article.
Table 3: Template of EN-Basic for each benchmark. #Test denotes the number of samples in the test set.
Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
high low high low high low high low high low high low high low
ChatGPT Native-Basic 44.4 19.4 84.6 69.7 56.9 48.6 51.6 40.6 35.1 36.4 32.5 29.9 50.8 40.8
EN-Basic 50.3 27.3 88.3 73.3 64.6 61.8 64.3 50.4 37.4 33.3 33.3 30.0 56.4 46.0
Native-CoT 65.1 27.1 84.1 69.8 54.9 47.4 51.6 43.4 35.5 35.1 31.9 27.9 53.8 41.8
EN-CoT 70.5 47.1 89.9 75.9 60.2 53.6 63.7 51.2 43.3 41.2 30.0 28.6 59.6 49.6
XLT 70.4 50.1 89.3 76.8 60.6 58.1 59.7 58.2 37.7 37.5 22.8 26.1 56.7 51.1
Trans-Google 74.7 72.7 90.3 83.2 62.4 59.1 68.2 62.0 42.5 48.3 30.6 28.9 61.4 59.0
Trans-NLLB 65.6 54.1 85.7 78.2 60.5 58.2 68.4 63.4 35.4 43.6 28.4 27.7 57.3 54.2
bloomz-7b1 Native-Basic 1.6 0.9 36.5 18.9 3.7 11.8 - - 7.1 10.5 - - 12.2 10.5
EN-Basic 1.9 2.2 67.5 55.2 48.2 40.7 - - 11.8 6.5 - - 32.4 26.2
Native-CoT 1.0 1.4 37.9 17.3 1.2 13.5 - - 5.2 11.1 - - 11.3 10.8
EN-CoT 1.7 1.6 61.3 52.8 37.6 34.7 - - 10.0 6.9 - - 27.7 24.0
XLT 1.9 1.5 58.6 49.2 35.4 35.3 - - 8.6 5.9 - - 26.1 23.0
Trans-Google 2.5 3.0 67.5 62.8 44.4 44.2 - - 15.6 23.0 - - 32.5 33.2
Trans-NLLB 2.0 2.9 64.3 61.2 44.1 43.6 - - 12.8 21.3 - - 30.8 32.2
Mistral-7B-Instruct Native-Basic 15.5 4.9 69.7 50.0 50.6 37.0 44.6 44.8 7.8 8.1 26.3 24.4 35.7 28.2
EN-Basic 33.7 8.8 42.5 33.8 55.5 46.2 47.0 46.6 6.8 8.0 21.7 21.1 34.5 27.4
Native-CoT 23.1 8.0 67.7 49.9 50.2 38.3 44.3 44.2 7.7 8.2 25.5 21.1 36.4 28.3
EN-CoT 37.3 13.1 50.9 38.9 54.2 46.8 46.6 46.4 11.3 12.0 18.7 18.8 36.5 29.3
XLT 43.0 15.0 78.3 57.9 48.4 44.3 47.9 47.2 9.4 10.4 17.1 19.6 40.7 32.4
Trans-Google 42.6 39.4 67.0 57.5 56.4 53.9 51.4 52.0 16.3 19.7 31.9 36.5 44.3 43.2
Trans-NLLB 32.3 30.8 62.1 52.3 54.4 51.9 52.2 53.6 14.5 19.3 31.0 37.3 41.1 40.9
Llama-2-13b-Chat Native-Basic 22.7 4.9 59.5 48.4 39.9 33.7 55.2 48.2 20.7 9.6 28.4 23.8 37.7 28.1
EN-Basic 28.7 4.4 63.9 51.6 48.2 39.8 59.6 56.8 20.9 17.8 31.3 30.2 42.1 33.4
Native-CoT 26.9 4.9 59.0 49.3 38.6 33.5 56.2 47.8 17.9 7.8 28.4 22.7 37.8 27.7
EN-CoT 29.5 5.5 68.2 51.0 46.2 41.8 57.8 56.6 20.5 17.3 30.7 28.0 42.1 33.4
XLT 32.8 6.5 68.1 52.7 56.9 47.3 56.0 54.2 19.6 16.8 22.0 18.1 42.6 32.6
Trans-Google 38.4 40.1 77.8 70.4 46.1 46.1 59.2 54.6 32.6 37.8 35.1 38.0 48.2 47.8
Trans-NLLB 32.8 30.4 72.7 67.1 45.6 45.2 58.1 56.2 26.7 34.7 33.4 37.3 44.9 45.1
Llama-2-70B-Chat Native-Basic 35.7 5.6 64.2 48.0 43.0 36.0 53.3 50.4 28.9 10.4 30.1 26.8 42.5 29.5
EN-Basic 42.5 7.7 70.7 52.0 52.7 41.9 61.9 52.8 25.7 21.5 30.2 35.3 47.3 35.2
Native-CoT 35.5 5.6 65.3 46.8 41.0 35.6 56.0 49.6 25.3 9.9 26.0 25.2 41.5 28.8
EN-CoT 45.6 7.0 80.7 56.3 52.7 40.9 66.5 57.0 32.7 25.7 29.8 32.0 51.3 36.5
XLT 49.0 8.4 76.4 54.7 57.3 48.4 56.6 51.6 26.5 26.7 19.3 11.5 47.5 33.6
Trans-Google 55.5 50.0 86.3 79.7 55.3 53.0 69.4 64.2 38.7 43.1 33.1 36.7 56.4 54.4
Trans-NLLB 46.5 39.7 83.3 75.6 53.7 51.0 70.5 62.4 17.8 24.7 32.4 36.2 50.7 48.3
Table 4: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting.
Model Prompt type de ru fr zh es ja sw th bn te avg
ChatGPT Native-Basic 48.8 42.8 42.8 36.0 50.0 46.0 30.8 21.6 15.6 9.6 34.4
EN-Basic 49.2 56.0 48.4 52.4 57.2 38.4 42.0 27.2 28.8 11.2 41.1
Native-CoT 66.0 69.6 62.4 64.4 70.0 58.0 49.2 28.4 20.8 10.0 49.9
EN-CoT 74.8 72.4 71.2 67.2 75.2 62.0 58.0 51.6 52.8 26.0 61.1
XLT 70.8 73.6 69.6 68.8 72.8 66.8 65.6 56.8 50.8 27.2 62.3
Trans-Google 76.8 76.4 75.2 73.2 76.0 70.4 73.6 76.0 74.0 67.2 73.9
Trans-NLLB 70.0 63.2 71.2 58.4 71.6 59.2 61.2 44.4 55.6 55.2 61.0
bloomz-7b1 Native-Basic 1.2 1.2 2.0 2.8 1.6 0.8 0.8 0.0 1.6 1.2 1.3
EN-Basic 2.0 1.6 2.4 2.8 1.6 1.2 2.0 1.2 3.6 2.0 2.0
Native-CoT 0.0 0.4 1.2 1.6 1.6 1.2 2.4 0.4 1.2 1.6 1.2
EN-CoT 2.0 1.2 2.4 2.0 0.8 2.0 1.6 1.2 2.0 1.6 1.7
XLT 0.8 1.2 2.0 3.2 1.6 2.4 2.0 0.8 0.8 2.4 1.7
Trans-Google 3.2 2.0 2.4 2.4 2.4 2.4 2.0 3.2 2.0 4.8 2.7
Trans-NLLB 2.4 1.6 3.2 0.8 2.0 2.0 3.6 2.4 2.4 3.2 2.4
Mistral-7B-Instruct Native-Basic 7.6 14.4 12.0 19.2 30.8 8.8 4.0 4.4 6.8 4.4 11.2
EN-Basic 38.4 36.4 31.6 28.0 42.4 25.6 7.6 9.6 16.0 2.0 23.8
Native-CoT 9.6 24.0 16.8 26.8 38.4 22.8 6.0 7.6 17.2 1.2 17.0
EN-CoT 39.2 42.0 36.0 33.6 42.0 30.8 8.0 21.6 18.4 4.4 27.6
XLT 43.6 51.6 45.2 38.4 45.2 34.0 10.4 23.6 19.6 6.4 31.8
Trans-Google 42.0 46.8 41.2 44.0 42.0 39.6 38.8 35.6 42.0 41.2 41.3
Trans-NLLB 37.6 30.0 34.0 24.8 38.0 29.6 31.6 26.4 31.2 34.0 31.7
Llama-2-13b-Chat Native-Basic 25.2 20.0 25.6 24.4 22.0 18.8 3.6 7.2 5.2 3.6 15.6
EN-Basic 32.4 26.4 32.0 26.0 34.8 20.8 3.2 5.6 5.6 3.2 19.0
Native-CoT 29.2 23.6 29.2 27.6 28.4 23.2 2.8 7.2 6.4 3.2 18.1
EN-CoT 34.0 32.4 32.0 24.4 35.6 18.4 5.6 6.8 6.0 3.6 19.9
XLT 34.4 34.4 33.6 29.6 37.2 27.6 4.8 8.4 9.2 3.6 22.3
Trans-Google 38.0 40.4 36.8 35.6 44.8 34.8 38.4 39.2 42.8 40.0 39.1
Trans-NLLB 29.6 33.2 38.8 31.2 28.0 36.0 32.0 24.8 35.6 29.2 31.8
Llama-2-70B-Chat Native-Basic 34.8 28.4 38.8 38.8 41.2 32.0 4.4 8.4 7.6 2.0 23.6
EN-Basic 50.4 39.2 48.0 40.0 48.0 29.6 6.0 8.8 11.6 4.4 28.6
Native-CoT 41.2 31.6 36.4 35.6 36.8 31.2 6.4 5.2 9.2 1.6 23.5
EN-CoT 49.6 48.0 50.0 38.0 48.4 39.6 7.6 7.2 10.4 2.8 30.2
XLT 52.0 49.6 49.6 47.2 52.0 43.6 8.0 8.0 15.6 2.0 32.8
Trans-Google 56.8 56.4 54.4 54.8 56.4 54.0 51.6 46.0 51.6 50.8 53.3
Trans-NLLB 49.6 43.6 49.2 41.2 50.4 45.2 43.6 32.0 42.0 41.2 43.8
Table 5: Accuracy scores across various languages on the MGSM benchmark.
Model Prompt type zh it vi tr id sw th et ta ht qu avg
ChatGPT Native-Basic 88.0 91.8 74.0 81.4 85.4 77.2 65.2 85.4 49.6 63.4 50.0 72.3
EN-Basic 90.0 89.8 85.0 86.0 87.2 78.2 75.0 81.4 58.2 65.8 54.8 76.1
Native-CoT 87.0 92.6 72.8 80.8 83.8 75.4 66.8 84.8 48.6 63.2 55.2 72.4
EN-CoT 90.4 92.2 87.0 89.6 90.2 85.6 74.8 85.8 61.4 69.2 50.2 78.6
XLT 89.4 91.2 87.4 88.0 88.8 82.4 76.4 91.0 60.6 76.8 50.4 79.3
Trans-Google 90.8 91.6 88.4 85.8 88.8 79.4 82.6 88.2 85.6 81.6 73.2 84.5
Trans-NLLB 85.6 89.2 82.4 85.8 87.0 81.4 73.8 85.4 80.6 76.2 55.6 79.7
bloomz-7b1 Native-Basic 46.6 48.6 14.4 1.6 48.4 39.0 20.0 0.0 19.0 2.8 20.6 21.4
EN-Basic 78.2 55.6 68.6 50.2 62.8 56.8 49.6 50.0 71.4 50.0 50.4 56.5
Native-CoT 43.4 50.0 20.2 0.6 48.6 23.0 39.2 0.0 17.6 0.0 9.4 20.9
EN-CoT 67.4 53.4 63.0 50.4 57.4 51.4 49.6 49.4 64.0 49.6 50.6 53.9
XLT 63.8 49.6 62.4 45.6 64.0 49.0 51.2 46.0 52.8 48.0 36.6 50.5
Trans-Google 68.0 68.6 66.0 65.2 68.8 60.4 59.4 67.2 61.8 61.6 57.6 63.7
Trans-NLLB 64.0 67.2 61.6 63.6 64.6 62.2 57.4 62.8 62.8 61.6 54.2 61.8
Mistral-7B-Instruct Native-Basic 67.2 82.2 59.8 55.0 65.0 47.6 51.8 36.6 49.2 51.2 43.6 54.2
EN-Basic 48.6 43.6 35.4 30.6 43.6 37.8 39.8 28.6 35.2 29.4 25.0 34.9
Native-CoT 64.0 80.4 58.6 54.6 65.4 45.4 50.0 40.0 44.2 51.2 48.2 53.8
EN-CoT 55.8 52.2 44.6 43.8 52.2 39.8 46.0 32.6 29.2 39.4 28.2 40.8
XLT 82.6 81.4 70.8 66.8 77.8 47.8 64.2 53.6 52.0 56.6 44.0 61.5
Trans-Google 69.4 64.8 66.8 61.0 68.8 52.2 62.0 60.8 59.8 52.0 43.6 59.2
Trans-NLLB 60.8 63.4 62.2 59.2 63.0 50.8 51.4 60.6 55.0 51.0 27.4 54.4
Llama-2-13b-Chat Native-Basic 65.0 62.2 51.4 50.4 57.6 46.2 48.4 50.0 40.2 47.2 47.0 50.1
EN-Basic 61.2 74.2 56.2 52.8 62.0 52.0 50.6 50.6 50.2 46.4 48.4 54.3
Native-CoT 62.8 64.6 49.6 53.8 64.8 49.8 51.8 45.4 32.6 49.8 46.6 50.9
EN-CoT 67.4 71.8 65.4 51.4 68.2 48.2 49.0 46.8 48.6 50.4 45.6 54.5
XLT 65.4 72.6 66.2 57.2 70.0 47.0 49.2 50.8 50.2 50.6 46.6 56.0
Trans-Google 77.8 80.4 75.2 75.0 76.4 66.6 67.6 74.0 71.8 68.8 63.2 71.9
Trans-NLLB 73.0 75.6 69.6 74.4 73.2 67.4 62.4 73.8 66.2 68.0 51.2 68.2
Llama-2-70B-Chat Native-Basic 61.6 81.6 49.4 49.4 55.4 50.6 46.8 49.8 41.0 46.4 44.6 51.5
EN-Basic 74.6 79.4 58.0 53.6 63.2 48.8 50.2 49.4 50.4 49.0 51.0 55.3
Native-CoT 65.8 78.0 52.2 51.8 54.8 49.2 49.2 50.2 40.0 43.2 36.2 50.5
EN-CoT 80.4 88.0 73.6 65.4 77.8 53.0 50.0 56.0 48.0 49.8 50.6 61.2
XLT 79.8 82.0 67.4 64.6 74.4 49.8 51.8 55.0 47.8 46.2 48.2 58.7
Trans-Google 87.2 88.0 83.6 82.2 89.4 76.6 77.6 83.4 83.6 76.4 68.4 80.9
Trans-NLLB 83.2 86.6 80.2 79.8 85.8 74.4 71.4 79.2 79.6 76.2 58.2 77.1
Table 6: Accuracy scores across various languages on the XCOPA benchmark.
Model Prompt type de ru fr zh es vi tr sw ar el th bg hi ur avg
ChatGPT Native-Basic 59.0 58.8 60.2 54.0 60.2 49.2 51.6 51.0 50.6 58.0 39.6 54.8 42.8 40.4 52.2
EN-Basic 68.6 58.2 67.4 62.2 68.4 63.0 65.6 65.2 62.4 64.6 56.4 65.4 55.8 59.0 63.0
Native-CoT 59.4 54.2 58.0 51.8 58.6 47.6 53.0 50.8 51.2 54.6 37.2 54.4 40.2 37.4 50.6
EN-CoT 62.6 56.4 61.4 57.4 65.8 57.6 58.0 54.0 53.4 59.0 51.0 59.8 48.6 45.0 56.4
XLT 63.0 57.8 61.4 58.4 63.4 59.8 61.4 58.0 57.8 60.4 55.0 59.6 53.2 59.2 59.2
Trans-Google 65.6 59.6 65.2 62.6 62.6 58.6 60.4 57.6 63.2 62.2 56.4 60.0 57.0 55.8 60.5
Trans-NLLB 63.4 62.2 61.6 57.4 62.8 55.6 59.4 58.8 62.4 63.4 54.2 61.6 52.8 53.0 59.2
bloomz-7b1 Native-Basic 0.4 13.4 0.2 6.6 1.4 0.0 6.8 18.2 1.6 5.2 26.6 15.4 17.8 2.8 8.3
EN-Basic 39.8 42.8 50.8 52.4 52.2 51.4 34.2 42.4 45.6 37.2 33.8 40.4 49.2 43.0 43.9
Native-CoT 0.4 3.0 1.2 1.2 1.2 0.2 9.0 27.2 1.6 0.8 33.4 12.4 20.0 3.8 8.2
EN-CoT 36.2 35.2 37.4 42.2 37.4 37.2 33.2 34.8 36.2 33.6 33.2 34.2 37.6 34.4 35.9
XLT 38.2 34.4 35.0 34.0 35.0 36.0 37.4 35.4 34.6 35.6 35.0 36.6 33.8 34.0 35.4
Trans-Google 45.0 43.4 44.2 44.0 45.2 44.8 43.8 44.0 44.0 44.6 44.4 44.8 43.4 44.4 44.3
Trans-NLLB 45.6 43.0 44.0 44.0 45.4 42.4 43.6 43.4 44.6 44.6 43.2 44.8 42.8 42.0 43.8
Mistral-7B-Instruct Native-Basic 50.4 55.6 59.2 46.0 59.0 33.4 38.8 33.0 34.2 34.2 39.2 46.6 37.0 33.2 42.8
EN-Basic 56.4 54.6 59.8 54.0 56.8 51.4 46.8 37.6 45.8 49.4 47.0 54.4 46.4 41.8 50.2
Native-CoT 50.0 55.0 58.4 47.6 54.6 35.8 38.2 32.2 37.6 35.4 40.0 52.0 36.8 33.8 43.4
EN-CoT 55.0 52.2 58.0 52.4 57.0 50.4 48.0 38.0 48.6 51.2 45.8 54.2 46.8 42.0 50.0
XLT 48.2 44.6 49.6 49.4 52.4 46.0 48.0 39.0 42.2 46.4 45.4 46.6 44.0 42.6 46.0
Trans-Google 58.6 54.2 59.2 52.6 59.0 55.0 54.6 53.0 56.4 58.2 48.8 56.8 52.4 50.6 55.0
Trans-NLLB 57.0 52.4 55.8 50.2 58.2 53.0 54.2 49.4 53.0 56.4 47.4 55.2 50.0 49.6 53.0
Llama-2-13b-Chat Native-Basic 41.4 40.2 44.0 38.6 42.8 32.4 34.6 31.6 32.8 34.2 34.0 37.4 31.4 33.6 36.4
EN-Basic 50.2 47.4 51.6 45.0 51.8 43.0 41.8 37.8 38.8 42.6 36.4 45.0 38.4 37.8 43.4
Native-CoT 39.4 42.0 43.4 32.6 42.6 31.8 31.4 33.4 31.2 35.2 32.8 38.2 32.6 33.2 35.7
EN-CoT 45.6 46.8 48.8 44.4 46.6 44.8 41.8 38.6 43.2 43.4 38.6 46.2 42.0 40.8 43.7
XLT 59.6 55.8 56.4 54.0 59.8 55.6 48.2 37.8 49.4 49.0 44.4 52.0 48.4 49.2 51.4
Trans-Google 50.4 44.2 45.4 44.6 46.0 46.0 47.6 42.8 48.4 48.2 43.4 45.4 45.4 47.4 46.1
Trans-NLLB 48.6 46.6 47.0 43.2 44.6 43.6 49.0 43.2 44.0 46.0 41.6 48.6 45.6 43.4 45.4
Llama-2-70B-Chat Native-Basic 44.0 42.0 45.4 42.6 45.6 38.4 38.4 32.6 35.0 37.6 33.0 41.8 34.8 34.8 39.0
EN-Basic 53.6 54.6 57.0 49.6 55.6 46.0 42.8 32.4 50.2 46.2 38.6 52.4 37.6 34.8 46.5
Native-CoT 40.4 42.2 45.4 38.4 41.4 38.4 36.6 32.8 35.2 37.4 32.6 41.0 33.2 36.2 37.9
EN-CoT 53.6 52.8 56.4 50.4 56.8 46.0 40.6 33.4 44.6 47.8 38.2 48.2 37.6 36.6 45.9
XLT 56.0 59.4 59.6 55.2 61.2 52.6 51.4 36.4 44.4 55.4 44.6 57.8 51.2 45.8 52.2
Trans-Google 58.8 53.4 56.8 56.4 54.8 51.8 55.4 49.6 57.2 56.4 50.2 57.4 50.8 46.6 54.0
Trans-NLLB 56.4 52.8 54.6 49.8 58.6 50.2 53.4 51.0 52.0 56.0 48.8 52.4 49.0 45.6 52.2
Table 7: Accuracy scores across various languages on the XNLI benchmark.
Model Prompt type de fr zh es ja ko avg
ChatGPT Native-Basic 62.0 53.6 46.6 46.6 49.0 40.6 49.7
EN-Basic 67.6 68.0 58.8 71.4 55.8 50.4 62.0
Native-CoT 61.8 55.0 48.6 48.8 44.0 43.4 50.3
EN-CoT 67.6 64.0 61.2 70.0 55.8 51.2 61.6
XLT 57.4 63.8 59.8 59.2 58.2 58.2 59.4
Trans-Google 69.0 69.6 66.0 71.4 65.0 62.0 67.2
Trans-NLLB 67.0 70.6 68.6 70.2 65.4 63.4 67.5
Mistral-7B-Instruct Native-Basic 40.6 47.0 49.2 44.2 41.8 44.8 44.6
EN-Basic 46.8 47.8 47.8 46.8 45.8 46.6 46.9
Native-CoT 43.8 50.2 38.8 43.6 45.0 44.2 44.3
EN-CoT 46.2 47.4 47.8 47.0 44.8 46.4 46.6
XLT 47.4 49.6 47.6 46.6 48.2 47.2 47.8
Trans-Google 51.2 49.8 54.0 49.6 52.4 52.0 51.5
Trans-NLLB 50.6 52.8 52.4 50.8 54.2 53.6 52.4
Llama-2-13b-Chat Native-Basic 50.8 57.2 54.0 58.0 55.8 48.2 54.0
EN-Basic 60.2 61.0 58.6 59.8 58.2 56.8 59.1
Native-CoT 50.4 58.8 59.0 55.8 56.8 47.8 54.8
EN-CoT 59.2 55.8 58.6 59.2 56.4 56.6 57.6
XLT 54.8 58.0 53.6 56.6 56.8 54.2 55.7
Trans-Google 56.6 62.0 59.6 61.6 56.2 54.6 58.4
Trans-NLLB 56.2 60.0 57.4 59.4 57.6 56.2 57.8
Llama-2-70B-Chat Native-Basic 53.4 49.8 55.6 61.0 46.8 50.4 52.8
EN-Basic 62.8 66.2 58.4 67.0 55.2 52.8 60.4
Native-CoT 53.0 53.4 53.6 65.4 54.6 49.6 54.9
EN-CoT 65.0 70.8 65.0 70.2 61.6 57.0 64.9
XLT 57.0 61.6 57.6 57.2 49.4 51.6 55.7
Trans-Google 70.6 70.6 68.0 72.2 65.6 64.2 68.5
Trans-NLLB 69.8 73.4 69.4 71.2 68.8 62.4 69.2
Table 8: Accuracy scores across various languages on the PAWS-X benchmark.
Model Prompt type de ru fr zh es ja vi tr th avg
ChatGPT Native-Basic 44.1 30.5 46.4 31.4 40.2 20.2 33.0 39.2 33.6 35.4
EN-Basic 36.9 30.5 43.3 28.9 44.1 43.5 34.7 32.7 34.0 36.5
Native-CoT 43.6 22.2 46.1 30.0 38.3 33.9 34.1 38.3 31.9 35.4
EN-CoT 44.6 37.4 49.7 38.5 48.0 52.4 32.6 42.0 40.5 42.9
XLT 36.6 31.0 39.3 31.8 44.0 43.6 37.3 37.9 37.2 37.6
ransg 42.0 39.2 42.7 48.6 40.8 46.4 37.8 44.2 52.3 43.8
Trans-NLLB 39.2 34.6 26.7 31.6 29.1 45.3 41.2 41.2 45.9 37.2
bloomz-7b1 Native-Basic 0.6 3.0 7.6 12.1 11.2 7.5 7.6 0.0 20.9 7.8
EN-Basic 7.5 3.7 12.3 21.4 12.2 12.3 13.3 2.1 11.0 10.6
Native-CoT 0.2 0.9 5.9 8.6 8.3 6.0 6.7 0.0 22.2 6.5
EN-CoT 4.0 3.0 11.4 17.9 13.9 8.7 11.1 1.7 12.2 9.3
XLT 5.7 2.8 10.2 14.8 10.1 7.1 9.6 1.4 10.4 8.0
Trans-Google 13.5 11.5 10.7 25.7 12.5 22.5 12.8 11.7 34.2 17.2
Trans-NLLB 11.7 8.7 7.2 15.2 9.3 24.5 13.1 11.2 31.3 14.7
Mistral-7B-Instruct Native-Basic 8.5 5.2 8.7 7.2 9.5 7.4 8.0 2.6 13.5 7.8
EN-Basic 7.9 5.0 7.5 5.1 8.7 6.7 6.3 5.3 10.6 7.0
Native-CoT 9.1 5.4 7.7 8.1 8.2 7.9 7.3 2.8 13.6 7.8
EN-CoT 11.2 7.8 16.0 8.4 14.9 13.1 7.9 7.6 16.4 11.5
XLT 9.7 7.2 10.4 8.4 10.4 10.5 9.2 6.6 14.2 9.6
Trans-Google 14.6 13.8 14.9 17.7 17.0 22.5 13.4 15.1 24.4 17.0
Trans-NLLB 13.3 12.7 10.5 14.9 11.8 24.1 13.8 13.5 25.2 15.5
Llama-2-13b-Chat Native-Basic 15.0 13.6 31.3 20.6 29.7 13.8 21.2 5.8 13.4 18.3
EN-Basic 28.5 11.6 28.7 13.9 27.2 21.0 15.3 15.6 20.0 20.2
Native-CoT 14.6 10.4 29.1 13.3 23.6 10.5 23.8 5.6 10.1 15.7
EN-CoT 28.2 12.6 31.1 11.9 28.9 15.3 15.4 18.3 16.3 19.8
XLT 23.6 17.0 27.5 10.3 26.2 18.2 14.7 16.4 17.2 19.0
Trans-Google 31.1 29.9 34.6 35.1 31.7 35.4 30.8 31.7 43.9 33.8
Trans-NLLB 26.1 26.6 19.8 27.4 18.5 36.2 32.1 29.2 40.2 28.4
Llama-2-70B-Chat Native-Basic 36.7 23.8 35.2 15.9 39.3 24.7 26.7 8.6 12.1 24.8
EN-Basic 33.2 18.1 32.9 18.8 33.7 26.6 16.3 20.7 22.3 24.7
Native-CoT 34.8 19.5 33.9 13.1 38.5 13.1 24.1 9.2 10.6 21.9
EN-CoT 39.5 24.6 39.0 24.2 41.0 35.2 25.3 26.4 25.0 31.1
XLT 29.8 22.4 29.6 18.0 31.3 29.5 25.0 27.3 26.1 26.6
Trans-Google 37.3 34.0 37.1 43.5 35.4 48.0 35.8 38.3 47.9 39.7
Trans-NLLB 16.7 16.4 11.9 18.5 14.9 26.8 19.7 21.8 27.6 19.4
Table 9: F1 scores across various languages on the MKQA benchmark.
Model Prompt type fr zh es vi tr avg
ChatGPT Native-Basic 29.2 39.3 26.9 34.4 29.9 31.9
EN-Basic 28.9 38.8 27.8 37.9 30.0 32.7
Native-CoT 28.8 38.5 26.1 34.0 27.9 31.1
EN-CoT 25.4 35.1 26.0 33.5 28.6 29.7
XLT 24.2 25.5 18.1 23.4 26.1 23.4
Trans-Google 27.2 36.2 26.3 32.6 28.9 30.3
Trans-NLLB 26.4 29.7 26.1 31.5 27.7 28.3
bloomz-7b1 Native-Basic 14.6 24.3 20.0 7.7 8.2 14.9
EN-Basic 20.1 23.9 20.9 20.6 14.2 19.9
Native-CoT 18.2 25.4 24.1 1.7 8.0 15.5
EN-CoT 18.0 26.1 21.6 19.3 11.3 19.3
XLT 12.2 19.9 19.3 14.5 5.3 14.2
Trans-Google 10.0 14.2 12.1 9.0 10.7 11.2
Trans-NLLB 10.5 8.6 12.5 9.7 11.5 10.6
Mistral-7B-Instruct Native-Basic 23.0 34.0 22.3 25.8 24.4 25.9
EN-Basic 20.9 16.5 21.5 28.0 21.1 21.6
Native-CoT 19.7 33.6 22.1 26.4 21.1 24.6
EN-CoT 20.6 12.1 19.9 22.2 18.8 18.7
XLT 15.4 16.5 14.7 21.7 19.6 17.6
Trans-Google 26.8 34.9 26.4 39.5 36.5 32.8
Trans-NLLB 26.8 30.0 26.6 40.6 37.3 32.2
Llama-2-13b-Chat Native-Basic 27.7 21.9 25.3 38.8 23.8 27.5
EN-Basic 25.7 38.2 23.6 37.7 30.2 31.1
Native-CoT 27.9 29.0 24.8 31.8 22.7 27.2
EN-CoT 24.0 39.4 23.1 36.4 28.0 30.2
XLT 24.2 17.7 22.4 23.6 18.1 21.2
Trans-Google 28.0 42.9 27.9 41.6 38.0 35.7
Trans-NLLB 27.5 37.5 26.9 41.6 37.3 34.2
Llama-2-70B-Chat Native-Basic 28.8 34.5 27.3 29.7 26.8 29.4
EN-Basic 29.0 31.8 24.3 35.7 35.3 31.2
Native-CoT 25.3 29.5 26.7 22.4 25.2 25.8
EN-CoT 27.0 35.2 22.1 34.8 32.0 30.2
XLT 18.1 29.7 15.2 14.2 11.5 17.7
Trans-Google 26.8 39.7 27.1 38.7 36.7 33.8
Trans-NLLB 26.6 37.5 26.3 39.0 36.2 33.1
Table 10: ROUGE-1 scores across various languages on the XL-sum benchmark.
Figure 6: The LLM-as-a-judge prompt for GPT-4-Turbo.
Figure 7: Prompt to check whether answering a request needs local cultural knowledge.